the number of homes built can be modeled by an exponential function,
H= p * at , where p is the number of new homes built in the first year recorded.
p=10 t =3 a=2
I'm not sure how to calculate : the number of homes built can be modeled by an exponential function,
H= p * a^t , where p is the number of new homes built in the first year recorded.
p=10 t =3 a=2
I'm not sure how to calculate

Answer by rapaljer(4551)

H= p*a^t
H = 10*2^3
H = 10*8
H = 80 homes

R^2